OpenStudy (anonymous):

Grammys apple orchard is selling apple cider in 34.7 ounce bottles for $2.09. Grammys competitor is selling 24.6 ounce bottles for $1.99. what is Grammys competitor price per ounce per cider.

OpenStudy (anonymous):

its almost the same question

OpenStudy (aric200):

?

OpenStudy (anonymous):

grammys competitor price poer ounce now

OpenStudy (aric200):

Ah per cider.

OpenStudy (aric200):

It says per cider not ounce..

OpenStudy (anonymous):

oh....u know what i mean...how come u nvr post questions

OpenStudy (aric200):

We already established that it is $0.06 Cents per ounce So if the bottle is 34.7 ounce bottles then Just Do 34.7 * 0.06

OpenStudy (aric200):

And I do post questions I have posted plenty most of the time they just go unanswered.

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

post one now i bet some one will answer it!

OpenStudy (aric200):

You get 2.082 But you need to round this off to the nearrest tenth so the answer would be $2.08

OpenStudy (anonymous):

it said it wrong

OpenStudy (anonymous):

brb

OpenStudy (aric200):

@mathmale Can you tell me what I did wrong?

OpenStudy (aric200):

I thought this was asking per bottle not ounce

OpenStudy (aric200):

But I still even said earlier " We already established that it is $0.06 Cents per ounce"

OpenStudy (johnweldon1993):

"what is Grammys COMPETITORS price per ounce per cider." "Grammys apple orchard is selling apple cider in 34.7 ounce bottles for $2.09" So for Grammy, we do \(\large \frac{2.09}{34.7} = .06cents\) per ounce "Grammys competitor is selling 24.6 ounce bottles for $1.99" And for her competitor we do \(\large \frac{1.99}{24.6} = .08cents\) per ounce

OpenStudy (anonymous):

back

OpenStudy (aric200):

I confussed myself haha

OpenStudy (anonymous):

how?

OpenStudy (johnweldon1993):

Yeah, you were correct in the.06 cents for grammys, however the question wants her competitors price

OpenStudy (anonymous):

yes

OpenStudy (aric200):

Yeah I was paying no attention forgive me

OpenStudy (anonymous):

i for gives u lml

OpenStudy (anonymous):

That's simply $2.09/34.7 = $0.06 per ounce ie 6 cents per ounce. The competitor is $1.99/24.6 = $0.08 per ounce ie 8 cents per ounce

OpenStudy (anonymous):

thanks for your 8 cents terrel...lol get it? instead of 2 cents i said 8 lol...ok that was corny

OpenStudy (anonymous):

lmao yea